the product of 4, 5, and 6 is equal to twice the sum of 10 and what number?
Answer by mgmoeab(37)

if understood right:
the product of 4 *5 *6 = 120
twice the sum of 10 and what number can be written as: 2 (10 + x) : where x is the number you are looking for
So.. if the product of 4, 5, and 6 is EQUAL to twice the sum of 10 and what number, then you solve like this:
120 = 2 ( 10 + x )
60 = 10 + x
x = 50
Verifying:
120 = 2 ( 10 + 50 )
120 = 2 ( 60 )
120 = 120